### v1

# Neutrino mass and ultralight dark matter mass from the Higgs mechanism

Abstract
We propose a model where the small neutrino masses are generated via the Higgs mechanism, mediated by an ultralight dark matter (ULDM) field with a mass $m\gtrsim 10^{-23}~{}eV$ and a characteristic energy scale of $O(10)~{}eV$. The resulting neutrino masses, as well as the mass and self-interaction strength of ULDM, are consistent with cosmological observations. The phase transition temperature for spontaneous symmetry breaking of ULDM occurs near the grand unification theory (GUT) scale. In this framework, neutrino masses can vary with spacetime, a feature that may be experimentally detectable by neutrino oscillation experiments. Additionally, we explore a scenario where the tiny ULDM mass is generated through radiative corrections via the Coleman-Weinberg mechanism, starting from a massless field theory. Our model addresses both the neutrino mass and ULDM mass puzzles through a unified mechanism, offering insights into possible extensions of the Standard Model.
I Introduction
Ultralight dark matter (ULDM), often referred to as fuzzy dark matter, ultralight axion, wave-$\psi$ or scalar field dark matter Baldeschi et al. (1983); Sin (1994); Lee and Koh (1996); Matos and Guzman (2000); Hu et al. (2000), is gaining attention as a plausible alternative to cold dark matter (CDM), as it retains the advantages of CDM on scales larger than galaxies while potentially solving the issues associated with cold dark matter at galactic scales (for a review see  Hui et al. (2017); Ferreira (2021); Rindler-Daller (2021); Lee (2018); Matos et al. (2024)). The ULDM model has been shown to address many mysteries of galaxies and black holes, such as small-scale issues Hu et al. (2000); Salucci et al. (2003); Navarro et al. (1996); de Blok et al. (2003); Tasitsiomi (2003), the typical size and mass of galaxies  Lee (2009), the satellite plane problem Park et al. (2022), and the final parsec problem Koo et al. (2024). The wave-like nature and characteristic length scale of ULDM have been recognized for their ability to address these issues.
The extremely small mass ($m\gtrsim 10^{-23}~{}eV$) of ULDM can be hypothesized to originate from the string axiverse Arvanitaki et al. (2010), or a weak breaking of some symmetry Sin (1994). For example, in the Coleman-Weinberg mechanism Coleman and Weinberg (1973), a very small mass can emerge as a classical symmetry is spontaneously broken through quantum corrections. The self-interactions of ULDM Lee and Koh (1996) required for a spontaneous symmetry breaking (SSB) may also be necessary to resolve tensions with certain observations in the free model (i.e., the fuzzy dark matter model) with $m\simeq 10^{-22}~{}eV$, such as the Lyman-alpha forest observations Iršič et al. (2017); Armengaud et al. (2017). In Lee and Koh (1996), it was proposed that including the self-interaction of ULDM drastically increases the length scale of the model and allows for a wider range of dark matter mass, which can help avoid the tensions.
The fact that neutrinos have mass ($m_{\nu}\lesssim 0.8~{}eV$)  Navas et al. (2024) implies the need to extend the Standard Model, as does the existence of dark matter. Therefore, it is natural to develop a model that connects these two mysteries. In particular, the extremely small mass of ultralight dark matter further suggests considering a connection with neutrinos, the lightest fermions in the Standard Model. In this sense, our approach in this work can be considered as the simplest extension of the Standard Model, addressing both the masses of neutrinos and ULDM.
In this paper, we propose that both the neutrino mass and the ULDM mass arise from a single mechanism: the Higgs mechanism  Higgs (1964) involving a gauge singlet or feasibly interacting dark Higgs scalar and its Yukawa coupling to a massless fermion. We also consider a specific example with the Coleman-Weinberg mechanism for SSB, starting with an initially massless Higgs-like scalar and a massless neutrino. In this scenario, the neutrino acquires mass through SSB, while ULDM gains mass through radiative corrections.
In Section II, we discuss a neutrino mass generation scenario via SSB and demonstrate that the mass and self-interaction of ULDM are consistent with cosmological observations. In Section III, we consider a Coleman-Weinberg mechanism for the massless scalar with a massless fermion. Finally, in Section IV, we discuss the results.
II Self-interacting Ultralight dark matter with a Broken Symmetry
The ULDM can be described by a complex scalar field $\phi$ with an effective action Lee and Koh (1996),
$S=\int\sqrt{-g}d^{4}x[\frac{-R}{16\pi G}-\frac{g^{\mu\nu}}{2}\phi^{*}_{;\mu}\phi_{;\nu}-V(\phi)],$ (1)
where the potential for SSB can be given by
$V(\phi)=-\frac{m^{2}}{2}|\phi|^{2}+\frac{\lambda|\phi|^{4}}{4}$ (2)
with a dimensionless self-coupling constant $\lambda$. This potential may have a $U(1)$ symmetry at high temperature. Note the minus sign of the quadratic term. The evolution of the field is described by the Klein-Gordon equation
$\square\phi+2\frac{dV}{d|\phi|^{2}}\phi=0,$
where $\square$ is the d’Alembertian. It is well known that after the symmetry is broken, the field attain a vacuum expectation value $\phi=v\simeq m/\sqrt{\lambda}$. One can expand the potential near $v$ and redefine the field around $v$ as $\phi-v\rightarrow\phi$ which has an effective mass $m_{\phi}=\sqrt{2}m$ and an effective coupling constant $\lambda_{\phi}=\lambda$. In the Newtonian limit, we can ignore cubic terms in galactic dynamics because odd-power terms like $|\phi|^{2}\phi\propto e^{-im_{\phi}t}$ oscillate too rapidly—at frequencies on the order of $m_{\phi}$—relative to galactic time scales.
From the equation of motion, one can see that the typical field value is approximately $\phi\sim m_{\phi}/\sqrt{\lambda}$, and therefore the typical energy density is $m_{\phi}^{2}\phi^{2}\sim m_{\phi}^{4}/\lambda$. It is useful to define the typical energy scale of the scalar field as
$\tilde{m}\equiv m_{\phi}/\lambda^{1/4},$ (3)
which turns out to be of order of $10~{}eV$ from many cosmological constraints (See below). In the context of thermal field theory, the SSB phase transition occurs at $T_{c}\simeq m_{\phi}/\sqrt{\lambda}=\tilde{m}/\lambda^{1/4}$, which introduces another relevant energy scale.
Many constraints on ULDM model come from the observation of galaxies which is usually nonrelativistic. Thus, it is useful to derive the Newtonian limit of the Klein-Gordon equation. For this we decompose the field as  Hui et al. (2017)
$\phi(t,\mathbf{x})=\frac{1}{\sqrt{2m_{\phi}}}\left[e^{-im_{\phi}t}\psi(t,\mathbf{x})+e^{im_{\phi}t}\psi^{*}(t,\mathbf{x})\right],$
where $\psi$ satisfies the following Schrödinger-Poisson equation  Lee and Koh (1996); Boehmer and Harko (2007); Chavanis (2011); Li et al. (2014)
$i\hbar\partial_{{t}}{\psi}$ $=$ $-\frac{\hbar^{2}}{2m_{\phi}}\nabla^{2}{\psi}+m_{\phi}{\Phi}{\psi}+\lambda|\psi|^{2}\psi,$ (4)
$\nabla^{2}{\Phi}$ $=$ ${4\pi G}\rho,$
where $\Phi$ is the gravitational potential, $\rho=m_{\phi}|\psi|^{2}$ is the dark matter density, and we ignore the expansion of the universe. In this paper we focus on the Thomas-Fermi regime where we can ignore quantum pressure compared to the pressure from the self-interaction. In this limit, the kinetic term in Eq. (4) can be ignored, and the exact ground state solution is given by  Lee and Koh (1996); Boehmer and Harko (2007); Chavanis (2011)
$|\psi|^{2}=\frac{|\psi(0)|^{2}}{\pi r/R_{TF}}sin\left(\frac{\pi r}{R_{TF}}\right),$ (5)
where the soliton size is
$R_{TF}={\sqrt{\frac{\pi\hbar^{3}\lambda}{8cGm_{\phi}^{4}}}}.$ (6)
This scale represents the typical size of a core of galaxies.
On the other hand, from the Schrödinger-Poisson equations one can obtain an equation for the density contrast $\delta\equiv\delta\rho/\bar{\rho}=\sum_{k}\delta_{k}e^{ik\cdot r}$ with a wave vector $k$ and the average matter density $\bar{\rho}$ of the universe,
$\frac{d^{2}\delta_{k}}{dt^{2}}+\left[(c^{2}_{q}+c^{2}_{s})k^{2}-4\pi G\bar{\rho}\right]\delta_{k}=0,$ (7)
where $c_{q}=\hbar k/2m_{\phi}$ is a quantum velocity and $c_{s}=\sqrt{{\hbar^{3}\lambda\bar{\rho}}/{2cm_{\phi}^{4}}}$ is the sound velocity from the self-interaction. By equating $c_{s}^{2}k^{2}$ with $4\pi G\bar{\rho}$ one can obtain the Jeans length which is the characteristic length over which the gravity dominates over the repulsion:
$\lambda_{J}=2\pi/k_{J}={\sqrt{\frac{\pi\hbar^{3}\lambda}{2cGm_{\phi}^{4}}}}=0.978~{}kpc\left(\frac{\tilde{m}}{10~{}eV}\right)^{-2}.$ (8)
Note that $\lambda_{J}$ is independent of $\bar{\rho}$ and is $2R_{TF}$. These scales determine the typical size of small galaxies. From the data of dSphs ($R_{TF}\simeq 1kpc$) one can find $\tilde{m}\simeq 7~{}eV$ Diez-Tejedor et al. (2014). There are other cosmological constraints. To ensure that the oscillation of ULDM field behaves as dark matter at and after the time of matter-radiation equality, the quartic term in the potential should be smaller than the energy density at that time. This condition leads to  Li et al. (2014); Boudon et al. (2022) $\tilde{m}\gtrsim 1~{}eV$. An observational constraint associated with cluster mergers provides an upper bound on the dark matter cross-section, $\sigma/m_{\phi}\lesssim$ $1\mathrm{~{}cm}^{2}/\mathrm{g}$ which gives the upper bound Desjacques et al. (2018) $\lambda\lesssim 10^{-12}\left({m_{\phi}}/{1{eV}}\right)^{\frac{3}{2}}$. Considering all these facts Arbey et al. (2002), we can infer that a physical phenomenon at the energy scale of $\tilde{m}\simeq 10~{}\text{eV}$ is related to ULDM, and within the Standard Model, neutrinos are the only particles that have a similar energy scale.
Figure 1: The phase transition temperature for ULDM SSB as a function of $m_{\phi}$ with various values of $\tilde{m}\equiv m_{\phi}/\lambda^{1/4}$.
As an example, for the fiducial mass $m=10^{-22}~{}eV$ and $\tilde{m}\simeq 10~{}eV$, we obtain $\lambda=10^{-92}$ and $v=\frac{m}{\sqrt{\lambda}}=10^{15}$ GeV. Interestingly, this energy scale is similar to that of GUT. In this case the phase transition temperature for SSB also becomes $T_{c}\simeq v\simeq 10^{15}GeV$. More generally, $T_{c}$ is given by
$T_{c}\simeq 10^{15}GeV\left(\frac{m_{\phi}}{10^{-22}}\right)\left(\frac{\lambda}{10^{-92}}\right)^{-1/2}=10^{15}GeV\left(\frac{10^{-22}~{}eV}{m_{\phi}}\right),$ (9)
where we have used the relation $\tilde{m}\simeq 10~{}eV$ in the last equation. From the cosmological observations of galaxies ULDM theory reproduces the GUT scale naturally. This temperature is much higher than the typical Bose-Einstein transition temperature for fuzzy dark matter Urena-Lopez (2009). For a more massive $m_{\phi}$, $T_{c}$ can be lower. In that case a Pecci-Quinn phase transition with $T_{c}\lesssim 10^{13}~{}GeV$ can be another possible phase transition Hui et al. (2017). (See Fig. 1.)
We will show how a scalar field in a $\phi^{4}$ theory with an energy scale $\tilde{m}\simeq 10~{}\text{eV}$ and a very small interaction constant $\lambda\simeq 10^{-92}$ can be related to neutrino masses when the field $\phi$ acquires a large vacuum expectation value $v$ around $10^{15}~{}\text{GeV}$. The Yukawa interaction in this context is given by:
$\mathcal{L}_{\text{Yukawa}}=-y\phi\bar{\nu}\nu$ (10)
where $y$ is the Yukawa coupling constant between the neutrino field $\nu$ and the scalar field. We ignored helicity since here we are interested in the magnitude of the mass. After SSB the neutrino mass $m_{\nu}=yv$ becomes
$m_{\nu}=0.1~{}eV\left(\frac{y}{10^{-25}}\right)\left(\frac{v}{10^{15}\mathrm{GeV}}\right),$ (11)
which implies the SSB giving mass to neutrionos can happen during the GUT phase transition, possibly with the inflation. In that case ULDM can be related to the GUT Higgs field. Conversely, one can obtain the Yukawa coupling $y$ as:
$y=\frac{m_{\nu}\sqrt{\lambda}}{m_{\phi}}=10^{-25}\left(\frac{m_{\nu}}{0.1~{}eV}\right)\left(\frac{m_{\phi}}{10^{-22}eV}\right)^{-1}\left(\frac{\lambda}{10^{-92}}\right)^{1/2}\simeq 10^{-25}\left(\frac{m_{\nu}}{0.1~{}eV}\right)\left(\frac{m_{\phi}}{10^{-22}eV}\right)$ (12)
for $\tilde{m}\simeq 10~{}eV$. Note that $y\simeq m_{\nu}\lambda^{1/4}/(10~{}eV)$ and the one-loop quantum correction from the Yukawa term is $O(y^{4})\simeq 10^{-8}\lambda\ll\lambda$, which does not destabilize the potential. Another constraint comes from the propagation of astrophysical neutrinos from distant sources  Brdar et al. (2018), which is
$y\lesssim 1.3\times 10^{-11}\left(\frac{m_{\nu}}{0.1~{}eV}\right)^{1/2}\left(\frac{m_{\phi}}{10^{-22}eV}\right)^{3/4}.$ (13)
We see that the mass and self-interaction of ULDM are consistent with cosmological observations and safely give neutrino masses observed with an appropriate $y$.
III Coleman-Weinberg Mechanism
In this section we investigate a specific example with the Coleman-Weinberg Mechanism for generating the mass of ULDM. The smallness of ULDM mass and neutrino mass implies a slight breaking of some symmetry due to a quantum correction. We start with a Lagrangian that includes a massless fermion $\psi$ and a massless complex scalar field $\phi$ with a quartic interaction:
$\mathcal{L}=\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi^{*})-V_{0}(\phi)+i\overline{\psi}\not{\partial}\psi-y\overline{\psi}\phi\psi+H.C.,$ (14)
where the tree level potential $V_{0}(\phi)=\lambda|\phi|^{4}/4$ and we ignore the gravity and the handness for simplicity. At one-loop level, quantum corrections generate an effective potential for the scalar field. Using the dimensional regularization, the effective potential $V_{\text{eff}}(\phi)$ can be computed as  Coleman and Weinberg (1973):
$V_{\text{eff}}(\phi)=\frac{\lambda}{4}|\phi|^{4}+\frac{3\lambda^{2}|\phi|^{4}}{64\pi^{2}}\left(\log\frac{3\lambda|\phi|^{2}}{\mu^{2}}-\frac{3}{2}\right)-\frac{y^{4}|\phi|^{4}}{16\pi^{2}}\left(\log\frac{y^{2}|\phi|^{2}}{\mu^{2}}-\frac{3}{2}\right),$ (15)
where $\mu$ is a renormalization scale. Note that the fermion contribution is negative and can destabilize the potential. To stabilize the potential $V_{\text{eff}}(\phi)$ we require $\lambda^{2}\gg y^{4}$, which later turns out to be satisfied for the neutrino mass range observed. From the radiative correction the scalar field can acquire a nonzero vacuum expectation value, $v$, which minimizes $V_{\text{eff}}$, i.e.,
$\left.\frac{dV_{\text{eff}}}{d\phi}\right|_{\phi=v}=v^{3}\left[\lambda+\frac{3\lambda^{2}}{16\pi^{2}}\left(\ln\frac{3\lambda v^{2}}{\mu^{2}}-1\right)-\frac{y^{4}}{4\pi^{2}}\left(\ln\frac{y^{2}v^{2}}{\mu^{2}}-1\right)\right]=0.$ (16)
Solving for $\lambda$ in terms of the other terms and substituting it into $V_{\text{eff}}$ simplifies the expression
$V_{\text{eff}}=\left[\frac{3\lambda^{2}}{64\pi^{2}}-\frac{y^{4}}{16\pi^{2}}\right]\left(ln(\phi^{2}/v^{2})-\frac{1}{2}\right)\phi^{4},$ (17)
which spontaneously breaks the symmetry of $V_{0}$.
The effective mass of the scalar field $m_{\phi}$ near the true vacuum is given by the second derivative of the effective potential at $\phi=v$:
$m_{\phi}^{2}\equiv\left.\frac{d^{2}V_{\text{eff}}}{d\phi^{2}}\right|_{\phi=v}=\frac{\left(3\,\lambda^{2}-4\,y^{4}\right)v^{2}}{8\,\pi^{2}}$ (18)
The scalar self-interaction coupling constant $\lambda_{\phi}$ can be extracted from the fourth derivative of $V_{\text{eff}}(\phi)$. At leading order, it is
$\lambda_{\phi}\equiv\frac{1}{6}\left.\frac{d^{4}V_{\text{eff}}}{d\phi^{4}}\right|_{\phi=v}=\frac{11\left(3\,\lambda^{2}-4\,y^{4}\right)}{48\,\pi^{2}}\simeq 0.069\lambda^{2}$ (19)
which is proportional to $\lambda^{2}$. The stability condition $\lambda_{\phi}>0$ leads to $\lambda^{2}\gg y^{4}$. Unless there is a miraculous cancellation, $y$ should be much smaller than $\lambda_{\phi}^{1/4}\gtrsim 10^{-23}$, which is satisfied for the neutrino mass $m_{\nu}=yv<1eV$ for $v\simeq 10^{15}GeV$ (see Eq. (12)). The effective energy scale $\tilde{m}$ now becomes
$\frac{m_{\phi}}{\lambda_{\phi}^{1/4}}=\left(\frac{3\left(3\lambda^{2}-4y^{4}\right)}{44\pi^{2}}\right)^{1/4}v\simeq 0.38\lambda^{1/2}v\simeq 10~{}eV,$ (20)
which gives a bound for the parameters $\lambda,v$ and $y$. For example, from Eq. (19), the condition $\lambda_{\phi}=(m/\tilde{m})^{4}>0$ leads to
$y\lesssim 3.3\times 10^{-23}\left(\frac{\tilde{m}}{10~{}eV}\right)^{-1}\left(\frac{m_{\phi}}{10^{-22}eV}\right),$ (21)
which gives a new bound for neutrino masses $m_{\nu}\lesssim 33~{}eV$ for $v\simeq 10^{15}GeV$.
IV Discussions
We proposed a model in which the masses of neutrinos are generated through the Higgs mechanism by ULDM. We showed that ultralight mass of dark matter with the GUT scale phase transition can be related to neutrino mass ($\sim 0.1~{}eV$). This consistency is nontrivial, considering the vast difference in energy scales (from $10^{-22}~{}eV$ to $10^{24}eV$) involved in the calculations. Since $m_{\phi}$ is larger than the current Hubble parameter, there is a possibility that the ULDM field oscillates with a frequency of approximately $m_{\phi}$. Future data from NANOgrav  Agazie et al. (2023) or LISA  Amaro-Seoane et al. (2017) will possibly confirm this oscillation with gravitational wave detectors. Furthermore, it could induce variations in neutrino masses, which may be detectable in astrophysical or ground based neutrino oscillation experiments Brdar et al. (2018). Our model addresses both the neutrino and ULDM masses in a unified framework, which may offer insights into possible extensions of the Standard Model.