### v2

# Neutrino mass and ultralight dark matter mass from the Higgs mechanism

Abstract
We propose a model in which small neutrino masses are generated via Yukawa coupling to a self-interacting ultralight dark matter (ULDM) field, treated as a pseudo-Nambu-Goldstone boson associated with a heavy Higgs-like field. ULDM has a mass $m\gtrsim 10^{-22}~{}\text{eV}$ and a characteristic energy scale $\tilde{m}\simeq 10~{}\text{eV}$. The resulting neutrino mass, as well as the mass and self-interaction strength of ULDM, align with cosmological observations. A quantum stability condition for an ULDM effective potential demands a small mass for neutrinos roughly bounded by $\tilde{m}$. The phase transition temperature for the Higgs mechanism can approach the grand unified theory (GUT) scale, potentially inducing the electroweak scale by reverting the type I seesaw mechanism for Majonara neutrinos. In this framework, neutrino masses can vary with spacetime, a feature that may be experimentally detectable through neutrino oscillation experiments. We also explore a scenario in which the tiny ULDM mass arises through radiative corrections via the Coleman-Weinberg mechanism, beginning from a massless field theory. Our model addresses both the neutrino mass and ULDM mass puzzles through a unified approach, providing insights into possible extensions of the Standard Model.
I Introduction
Ultralight dark matter (ULDM), often referred to as fuzzy dark matter, ultralight axion, wave-$\psi$ or scalar field dark matter Baldeschi et al. (1983); Sin (1994); Lee and Koh (1996); Matos and Guzman (2000); Hu et al. (2000); Boehmer and Harko (2007), is gaining attention as a plausible alternative to cold dark matter (CDM), as it retains the advantages of CDM on scales larger than galaxies while potentially solving the issues associated with cold dark matter at galactic scales (for a review see  Hui et al. (2017); Ferreira (2021); Rindler-Daller (2021); Lee (2018); Matos et al. (2024)). The ULDM model has been shown to address many mysteries of galaxies and black holes, such as small-scale issues Hu et al. (2000); Salucci et al. (2003); Navarro et al. (1996); de Blok et al. (2003); Tasitsiomi (2003), the typical size and mass of galaxies  Lee (2009), the satellite plane problem Park et al. (2022), and the final parsec problem Koo et al. (2024); Bromley et al. (2024). The wave-like nature and a characteristic length scale about $kpc$ of ULDM have been recognized for their ability to address these issues.
The extremely small mass ($m\gtrsim 10^{-22}~{}eV$) of ULDM can be hypothesized to originate from a weak breaking of some symmetry Sin (1994) or the string axiverse Arvanitaki et al. (2010). For example, in the Coleman-Weinberg mechanism Coleman and Weinberg (1973), a very small mass can emerge as a classical symmetry is spontaneously broken through quantum corrections. In Ref. Sin, 1994, ULDM was proposed as a pseudo-Nambu-Goldstone boson (pNGB). The self-interactions of ULDM Lee and Koh (1996) for a spontaneous symmetry breaking (SSB) could also resolve the tensions with certain observations in the free model (i.e., the fuzzy dark matter model) with $m\simeq 10^{-22}~{}eV$, such as the Lyman-alpha forest observations Iršič et al. (2017); Armengaud et al. (2017). In Lee and Koh (1996), it was proposed that including the self-interaction of ULDM drastically increases the length scale of the model and allows for a wider range of dark matter mass, which can help avoid the tensions of the fuzzy dark matter.
The fact that neutrinos have mass ($m_{\nu}\lesssim 0.8~{}eV$)  Navas et al. (2024) implies the need to extend the Standard Model, as does the existence of dark matter. Therefore, it is natural to develop a model that connects these two mysteries. In particular, the extremely small mass of ULDM suggests considering a connection with neutrinos, the lightest fermions in the Standard Model. In this sense, our approach in this work can be considered as the simplest extension of the Standard Model, addressing both the masses of neutrinos and ULDM. There are proposals suggesting that ULDM can interact with neutrinos  Berlin (2016); Krnjaic et al. (2018); Brdar et al. (2018); Davoudiasl et al. (2018). Our work proposes SSB and a phase transition related to neutrino mass generation by ULDM, constrained by cosmological observations.
In this paper, we propose that the neutrino mass arises from the Higgs mechanism Higgs (1964) involving a standard model gauge singlet or a feasibly interacting dark scalar (ULDM) and its Yukawa coupling to a massless fermion. We assume ULDM to be a pNGB associated with a heavy Higgs-like field. In the first scenario, ULDM has mass and self-interaction at a tree level. We also consider another scenario with the Coleman-Weinberg mechanism for SSB, starting with an initially massless dark scalar and a massless neutrino. In the latter scenario, the neutrino acquires mass through SSB, while ULDM gains mass through radiative corrections explaining the small mass.
In Section II, we discuss a neutrino mass generation scenario via SSB at a tree-level and demonstrate that the mass and self-interaction of ULDM are consistent with cosmological observations. In Section III, we consider another scenario based on the one-loop Coleman-Weinberg mechanism for the massless scalar with a massless fermion. Finally, in Section IV, we discuss the results.
II Self-interacting Ultralight dark matter with a Broken Symmetry
The ULDM as a pNGB can be described by a real scalar field $\phi$ with an effective action Lee and Koh (1996); Chavanis (2011),
$S=\int\sqrt{-g}d^{4}x[\frac{-R}{16\pi G}-\frac{g^{\mu\nu}}{2}\phi_{;\mu}\phi_{;\nu}-V(\phi)],$ (1)
where the potential for SSB can be given by
$V(\phi)=-\frac{m^{2}}{2}\phi^{2}+\frac{\lambda\phi^{4}}{4}$ (2)
with a dimensionless self-coupling constant $\lambda$. Note the minus sign of the quadratic term. The evolution of the field is described by the Klein-Gordon equation
$\square\phi+2\frac{dV}{d\phi^{2}}\phi=0,$
where $\square$ is the d’Alembertian. It is well known that after the symmetry is broken, the field attain a vacuum expectation value
$\langle\phi\rangle=v\simeq m/\sqrt{\lambda}.$ (3)
One can expand the potential near $v$ and redefine the field around $v$ as $\phi-v\rightarrow\phi$ which has an effective mass $m_{\phi}=\sqrt{2}m$ and an effective coupling constant $\lambda_{\phi}=\lambda$. In the Newtonian limit, we can ignore cubic terms in galactic dynamics because odd-power terms like $\phi^{3}$ oscillate rapidly at frequencies on the order of $m_{\phi}$ and average out to zero relative to galactic time scales.
From the equation of motion, one can see that the typical field value is approximately $\phi\sim m_{\phi}/\sqrt{\lambda}$, and therefore the typical energy density is $m_{\phi}^{2}\phi^{2}\sim m_{\phi}^{4}/\lambda$. It is useful to define the typical energy scale of the scalar field as
$\tilde{m}\equiv m_{\phi}/\lambda^{1/4},$ (4)
which turns out to be of order of $10~{}eV$ from many cosmological constraints (See below of Eq. (11)). In the context of thermal field theory, the SSB phase transition occurs at
$T_{c}\simeq\alpha v\simeq\alpha m_{\phi}/\sqrt{\lambda}=\alpha\tilde{m}^{2}/m_{\phi},$ (5)
where the constant $\alpha$ should be $O(1)$, unless there is fine-tuning. Note that if $\phi$ itself is a Higgs-like field, $T_{c}$ should be approximately $m_{\phi}$ not about $m_{\phi}/\sqrt{\lambda}$. Instead, here we have assumed that $\phi$ is an angular component (similar to an axion) of a Higgs-like field $\Phi$ having SSB at $T_{c}$. That is,
$\Phi(x,t)=\left(\langle\Phi\rangle+\Phi_{r}(x,t)\right)e^{i\phi(x,t)/\langle\Phi\rangle},$ (6)
where $\langle\Phi\rangle$ is the vacuum expectation value of $\Phi$ with a some appropriate potential, and $\Phi_{r}$ is a radial part. Since $\langle\phi\rangle=v=O(|\langle\Phi\rangle|)$ in general, we can expect $T_{c}=O(v)$. Therefore, ULDM can be a pNGB boson associated with a phase transition of $\Phi$ at $T_{c}$. Note that both $\Phi$ and $\phi$ undergo their own SSB. This scenario also explains the small mass $m$ of ULDM as a pNGB. If ULDM is a pNGB, it could be a real field with a soft symmetry breaking potential $V(\phi)$ in Eq. (2), which gives $\langle\phi\rangle=m/\sqrt{\lambda}$ in the $angular$ direction of $\Phi$. In this paper we do not consider a cosine potential which gives an effective attractive quartic term ($\lambda<0$).
Many constraints on ULDM model come from the observation of galaxies which is usually nonrelativistic. Thus, it is useful to derive the Newtonian limit of the Klein-Gordon equation. For this we decompose the real field as  Hui et al. (2017)
$\phi(t,\mathbf{x})=\frac{1}{\sqrt{2m_{\phi}}}\left[e^{-im_{\phi}t}\psi(t,\mathbf{x})+e^{im_{\phi}t}\psi^{*}(t,\mathbf{x})\right],$
where $\psi$ satisfies the following Schrödinger-Poisson equation  Lee and Koh (1996); Boehmer and Harko (2007); Chavanis (2011); Li et al. (2014)
$i\hbar\partial_{{t}}{\psi}$ $=$ $-\frac{\hbar^{2}}{2m_{\phi}}\nabla^{2}{\psi}+m_{\phi}{U}{\psi}+\frac{\lambda|\psi|^{2}\psi}{2m_{\phi}^{2}},$ (7)
$\nabla^{2}{U}$ $=$ ${4\pi G}\rho,$
where $U$ is the gravitational potential, $\rho=m_{\phi}|\psi|^{2}$ is the dark matter density, and we ignore the expansion of the universe. $\psi$ represents a zero-temperature macroscopic wave function for condensed ULDM. In this paper we focus on the Thomas-Fermi regime where we can ignore quantum pressure compared to the pressure from the self-interaction. In this limit, the kinetic term in Eq. (7) can be ignored, and the exact ground state solution is given by  Lee and Koh (1996); Boehmer and Harko (2007); Chavanis (2011)
$|\psi|^{2}=\frac{|\psi(0)|^{2}}{\pi r/R_{TF}}sin\left(\frac{\pi r}{R_{TF}}\right),$ (8)
where the soliton size is
$R_{TF}={\sqrt{\frac{\pi\hbar^{3}\lambda}{8cGm_{\phi}^{4}}}}.$ (9)
This scale represents the typical size of a core of galaxies.
On the other hand, from the Schrödinger-Poisson equations one can obtain an equation for the density contrast $\delta\equiv\delta\rho/\bar{\rho}=\sum_{k}\delta_{k}e^{ik\cdot r}$ with a wave vector $k$ and the average matter density $\bar{\rho}$ of the universe,
$\frac{d^{2}\delta_{k}}{dt^{2}}+\left[(c^{2}_{q}+c^{2}_{s})k^{2}-4\pi G\bar{\rho}\right]\delta_{k}=0,$ (10)
where $c_{q}=\hbar k/2m_{\phi}$ is a quantum velocity and $c_{s}=\sqrt{{\hbar^{3}\lambda\bar{\rho}}/{2cm_{\phi}^{4}}}$ is a sound velocity from the self-interaction. By equating $c_{s}^{2}k^{2}$ with $4\pi G\bar{\rho}$ one can obtain the Jeans length Khlopov et al. (1985); Chavanis (2011) which is the characteristic length over which the gravity dominates over the repulsion:
$\lambda_{J}=2\pi/k_{J}={\sqrt{\frac{\pi\hbar^{3}\lambda}{2cGm_{\phi}^{4}}}}=0.978~{}kpc\left(\frac{\tilde{m}}{10~{}eV}\right)^{-2}.$ (11)
Note that $\lambda_{J}$ is independent of $\bar{\rho}$ and is $2R_{TF}$. These scales determine the typical size of small galaxies. From the data of dSphs ($R_{TF}\simeq 1kpc$) one can find $\tilde{m}\simeq 7~{}eV$ Diez-Tejedor et al. (2014). There are other cosmological constraints. To ensure that the oscillation of ULDM field behaves as dark matter at and after the time of matter-radiation equality, the quartic term in the potential should be smaller than the energy density at that time. This condition leads to  Li et al. (2014); Boudon et al. (2022) $\tilde{m}\gtrsim 1~{}eV$. An observational constraint associated with cluster mergers provides an upper bound on the dark matter cross-section, $\sigma/m_{\phi}\lesssim$ $1\mathrm{~{}cm}^{2}/\mathrm{g}$ which gives the upper bound Desjacques et al. (2018) $\lambda\lesssim 10^{-12}\left({m_{\phi}}/{1{eV}}\right)^{\frac{3}{2}}$. Considering all these facts Arbey et al. (2002), we can infer that a physical phenomenon at the energy scale of $\tilde{m}\simeq 10~{}\text{eV}$ is related to ULDM, and neutrinos are the only known particles that have a similar mass scale.
Figure 1: The phase transition temperature $T_{c}=\alpha v$ for $\alpha=1$, yielding pNGB ULDM $\phi$, is given as a function of $m_{\phi}$ with various values of $\tilde{m}\equiv m_{\phi}/\lambda^{1/4}$.
As an example, for the fiducial mass $m_{\phi}=10^{-22}~{}eV$ and $\tilde{m}\simeq 10~{}eV$, we obtain $\lambda=10^{-92}$ and $v=\frac{m_{\phi}}{\sqrt{\lambda}}=10^{15}$ GeV. Interestingly, this energy scale is similar to that of GUT. In this case the phase transition temperature for SSB also becomes $T_{c}\simeq\alpha v\simeq 10^{15}\alpha~{}GeV$. More generally, $T_{c}$ is given by
$T_{c}\simeq 10^{15}\alpha~{}GeV\left(\frac{m_{\phi}}{10^{-22}~{}eV}\right)\left(\frac{\lambda}{10^{-92}}\right)^{-1/2}=10^{15}\alpha~{}GeV\left(\frac{10^{-22}~{}eV}{m_{\phi}}\right),$ (12)
where we have used the relation $\tilde{m}\simeq 10~{}eV$ in the last equation. Therefore, from the observations of galaxies, the self-interacting ULDM theory reproduces the GUT scale naturally. This temperature is much higher than the typical Bose-Einstein transition temperature for fuzzy dark matter Urena-Lopez (2009). The GUT scale field value of ULDM can also explain the dark matter density of the present universe  Hui et al. (2017). For a more massive $m_{\phi}$, $T_{c}$ can be lower. In that case a Pecci-Quinn phase transition with $T_{c}\lesssim 10^{13}~{}GeV$ can be another possible phase transition Hui et al. (2017). (See Fig. 1.)
We will show how a scalar field in a $\phi^{4}$ theory with an energy scale $\tilde{m}\simeq 10~{}\text{eV}$ and a very small interaction constant $\lambda\simeq 10^{-92}$ can be related to neutrino masses when the field $\phi$ acquires a large vacuum expectation value $v$ around $10^{15}~{}\text{GeV}$. The Yukawa interaction in this context is given by:
$\mathcal{L}_{\text{Yukawa}}=-y\phi\bar{\nu^{c}}\nu$ (13)
where $y$ is the Yukawa coupling constant between the neutrino field $\nu$ and the scalar field. Since $\phi$ is a real field, $\nu$ should be a Majonara neutrino. After SSB the neutrino mass $m_{\nu}=yv$ becomes
$m_{\nu}=0.1~{}eV\left(\frac{y}{10^{-25}}\right)\left(\frac{v}{10^{15}\mathrm{GeV}}\right),$ (14)
which implies the SSB giving mass to neutrionos can happen during the GUT phase transition, possibly with the inflation. In that case ULDM can be related to the GUT Higgs field $\Phi$. One can obtain the Yukawa coupling $y$ as:
$y=\frac{m_{\nu}\sqrt{\lambda}}{m_{\phi}}=10^{-25}\left(\frac{m_{\nu}}{0.1~{}eV}\right)\left(\frac{m_{\phi}}{10^{-22}eV}\right)^{-1}\left(\frac{\lambda}{10^{-92}}\right)^{1/2}\simeq 10^{-25}\left(\frac{m_{\nu}}{0.1~{}eV}\right)\left(\frac{m_{\phi}}{10^{-22}eV}\right)$ (15)
for $\tilde{m}\simeq 10~{}eV$. Note that $y\simeq m_{\nu}\lambda^{1/4}/(10~{}eV)$ and the one-loop quantum correction from the Yukawa term is $O(y^{4})\simeq O(10^{-8})\lambda\ll\lambda$, which does not destabilize the potential. Since the correction is negative, a repulsive self-interaction coupling ($\lambda>0$) at the tree level is required for stability. Conversely, if we require the one-loop correction from the Yukawa term $(O(y^{4}))$ to be smaller than $\lambda$, we obtain
$m_{\nu}=y\frac{m_{\phi}}{\sqrt{\lambda}}\lesssim\lambda^{1/4}\frac{m_{\phi}}{\sqrt{\lambda}}=\tilde{m}.$
Thus, the stability against quantum corrections demands a small neutrino mass, $m_{\nu}\lesssim O(10)~{}\text{eV}$, which is bounded by the energy scale $\tilde{m}$ of ULDM. Another constraint comes from the propagation of astrophysical neutrinos from distant sources  Brdar et al. (2018), which is
$y\lesssim 1.3\times 10^{-11}\left(\frac{m_{\nu}}{0.1~{}eV}\right)^{1/2}\left(\frac{m_{\phi}}{10^{-22}eV}\right)^{3/4}.$ (16)
We see that the mass and self-interaction of ULDM are consistent with cosmological observations and safely give neutrino mass scales observed with an appropriate $y$.
Another interesting observation is that a combination of $T_{c}$ and $\tilde{m}$ gives an energy scale (for $\alpha=1$)
$T_{EW}\equiv\sqrt{\tilde{m}T_{c}}\simeq 3.1\times 10^{3}GeV\left(\frac{m_{\phi}}{10^{-22}~{}eV}\right)\left(\frac{\lambda}{10^{-92}}\right)^{-3/8},$ (17)
which is similar to the electroweak scale.This implies that the phase transition of ULDM may provide a hint to resolving the hierarchy problem as well. For a specific example, one can consider the following Lagrangian in a mass basis
[…]
where $M_{R}=O(T_{c})$ is a GUT scale mass for a heavy neutrino. By reverting the logic of the type I seesaw mechanism with heavy right handed neutrinos $\nu_{R}$, one can approximately transform the Lagrangian in a flavor basis as
[…]
where $\tilde{T}_{EW}\equiv\sqrt{yvM_{R}}\lesssim\sqrt{\tilde{m}T_{c}}$. Unlike the usual type I seesaw scenario, the left-handed neutrino acquires a mass long before the electroweak phase transition and can induce a new scale $\tilde{T}_{EW}=O(TeV)$ after the GUT-scale phase transition in the thermal history of the universe at $T_{c}\simeq 10^{15}GeV$.
III Coleman-Weinberg Mechanism
In this section we investigate another example with the one-loop Coleman-Weinberg mechanism for generating the mass of ULDM. The smallness of the ULDM mass and neutrino mass implies a slight breaking of some symmetry due to a quantum correction. We start with a Lagrangian that includes a massless fermion $\psi$ and a massless real scalar field $\phi$ with a quartic interaction:
$\mathcal{L}=\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi)-V_{0}(\phi)+i\overline{\psi}\not{\partial}\psi-y\overline{\psi^{c}}\phi\psi+h.c.,$ (20)
where the tree level potential $V_{0}(\phi)=\lambda\phi^{4}/4$ and we ignore the gravity for simplicity. At the one-loop level, quantum corrections generate an effective potential for the scalar field. Using the dimensional regularization, the effective potential $V_{\text{eff}}(\phi)$ can be computed as  Coleman and Weinberg (1973):
$V_{\text{eff}}(\phi)=\frac{\lambda}{4}\phi^{4}+\frac{3\lambda^{2}\phi^{4}}{64\pi^{2}}\left(\log\frac{3\lambda\phi^{2}}{\mu^{2}}-\frac{3}{2}\right)-\frac{y^{4}\phi^{4}}{16\pi^{2}}\left(\log\frac{y^{2}\phi^{2}}{\mu^{2}}-\frac{3}{2}\right),$ (21)
where $\mu$ is a renormalization scale. Note that the fermion contribution is negative. From the radiative correction the scalar field can acquire a nonzero vacuum expectation value, $v$, which minimizes $V_{\text{eff}}$, i.e.,
$\left.\frac{dV_{\text{eff}}}{d\phi}\right|_{\phi=v}=v^{3}\left[\lambda+\frac{3\lambda^{2}}{16\pi^{2}}\left(\ln\frac{3\lambda v^{2}}{\mu^{2}}-1\right)-\frac{y^{4}}{4\pi^{2}}\left(\ln\frac{y^{2}v^{2}}{\mu^{2}}-1\right)\right]=0.$ (22)
Solving for $\lambda$ in terms of the other terms and substituting it into $V_{\text{eff}}$ simplifies the expression
$V_{\text{eff}}=\left[\frac{3\lambda^{2}}{64\pi^{2}}-\frac{y^{4}}{16\pi^{2}}\right]\left(ln(\phi^{2}/v^{2})-\frac{1}{2}\right)\phi^{4},$ (23)
which spontaneously breaks the symmetry of $V_{0}$. To stabilize the potential $V_{\text{eff}}(\phi)$ we require $\lambda^{2}\gg y^{4}$, which turns out below to be satisfied for the neutrino mass range observed.
The effective mass of the scalar field $m_{\phi}$ near the true vacuum is given by the second derivative of the effective potential at $\phi=v$:
$m_{\phi}^{2}\equiv\left.\frac{d^{2}V_{\text{eff}}}{d\phi^{2}}\right|_{\phi=v}=\frac{\left(3\,\lambda^{2}-4\,y^{4}\right)v^{2}}{8\,\pi^{2}}.$ (24)
The scalar self-interaction coupling constant $\lambda_{\phi}$ can be extracted from the fourth derivative of $V_{\text{eff}}(\phi)$. At the leading order, it is
$\lambda_{\phi}\equiv\frac{1}{6}\left.\frac{d^{4}V_{\text{eff}}}{d\phi^{4}}\right|_{\phi=v}=\frac{11\left(3\,\lambda^{2}-4\,y^{4}\right)}{48\,\pi^{2}}=\frac{11}{6}\frac{m^{2}_{\phi}}{v^{2}}\simeq 0.069\lambda^{2}$ (25)
which is proportional to $\lambda^{2}$ when $\lambda^{2}\gg y^{4}$. Compared to the quartic term, higher order terms are negligible as far as $|\phi-v|\ll v$.
The stability condition $\lambda_{\phi}>0$ leads to $\lambda^{2}\gg y^{4}$. Unless there is a miraculous cancellation, $y$ should be much smaller than $\lambda_{\phi}^{1/4}\gtrsim 10^{-23}$, which is satisfied for the neutrino mass $m_{\nu}=yv<1eV$ for $v\simeq 10^{15}GeV$ (see Eq. (15)). The effective energy scale $\tilde{m}$ now becomes
$\frac{m_{\phi}}{\lambda_{\phi}^{1/4}}=\left(\frac{3\left(3\lambda^{2}-4y^{4}\right)}{44\pi^{2}}\right)^{1/4}v\simeq 0.38\lambda^{1/2}v\simeq 10~{}eV,$ (26)
which gives a bound for the parameters $\lambda,v$ and $y$ in turn. For example, from Eq. (25), the stability condition $\lambda_{\phi}=(m_{\phi}/\tilde{m})^{4}>0$ leads to
$y\lesssim 3.3\times 10^{-23}\left(\frac{\tilde{m}}{10~{}eV}\right)^{-1}\left(\frac{m_{\phi}}{10^{-22}eV}\right),$ (27)
which gives a new bound for neutrino masses $m_{\nu}\lesssim 33~{}eV$ for $v\simeq 10^{15}GeV$. This mass bound is also satisfied for light sterile neutrinos Feng et al. (2022).
IV Discussions
We proposed a model in which the masses of neutrinos are generated through the Higgs mechanism by pNGB ULDM. Since the one-loop correction of the Yukawa coupling induces a negative quartic coupling, a positive self-interaction coupling of ULDM is inevitable. We showed that ultralight mass of dark matter with the GUT scale phase transition temperature can be related to neutrino mass ($\sim 0.1~{}eV$) and the electroweak scale. This consistency is nontrivial, considering the vast difference in energy scales (from $10^{-22}~{}eV$ to $10^{24}eV$) involved in the calculations.
The challenges of the neutrino mass and the hierarchy problem can be reframed in terms of the small mass and self-interactions of ULDM, thereby reducing the overall number of fine-tuning issues in cosmology and particle physics. If ULDM is a pNGB, the small mass of ULDM particles arises naturally. Since $m_{\phi}$ is larger than the current Hubble parameter, there is a possibility that the ULDM field oscillates with a frequency of approximately $m_{\phi}$. The oscillation of ULDM could be described by a plane wave $\phi\simeq\sqrt{\frac{2\rho_{\odot}}{m_{\phi}}}\cos\left[m_{\phi}(t-\textbf{v}\cdot x)\right]$, where $\rho_{\odot}$ is the dark matter density near the sun Berlin (2016) and v is the velocity of the wave. Future data from NANOgrav  Agazie et al. (2023); Chowdhury et al. (2024) or LISA  Amaro-Seoane et al. (2017) will possibly confirm this oscillation with gravitational wave detectors. The oscillation of the field could induce variations in neutrino masses, which may be detectable in astrophysical or ground based neutrino oscillation experiments Berlin (2016); Krnjaic et al. (2018); Brdar et al. (2018) in the near future with an effective Langrangian $\mathcal{L}_{\text{eff}}=-m_{\nu}\left(1+\frac{\phi}{v}\right)\bar{\nu^{c}}\nu.$ Our model addresses both the neutrino and ULDM masses in a unified framework, which may offer insights into possible extensions of the Standard Model.
IV.1 acknowledgments
The author is thankful to Sin Kyu Kang for helpful comments.